# MS-Temba : Multi-Scale Temporal Mamba for Efficient Temporal Action Detection

### v1

@@ Abstract
Action detection in real-world scenarios is particularly challenging due to densely distributed actions in hour-long untrimmed videos. It requires modeling both short- and long-term temporal relationships while handling significant intra-class temporal variations. Previous state-of-the-art (SOTA) Transformer-based architectures, though effective, are impractical for real-world deployment due to their high parameter count, GPU memory usage, and limited throughput, making them unsuitable for very long videos. In this work, we innovatively adapt the Mamba architecture for action detection and propose Multi-scale Temporal Mamba (MS-Temba), comprising two key components: Temporal Mamba (Temba) Blocks and the Temporal Mamba Fuser. Temba Blocks include the Temporal Local Module (TLM) for short-range temporal modeling and the Dilated Temporal SSM (DTS) for long-range dependencies. By introducing dilations, a novel concept for Mamba, TLM and DTS capture local and global features at multiple scales. The Temba Fuser aggregates these scale-specific features using Mamba to learn comprehensive multi-scale representations of untrimmed videos. MS-Temba is validated on three public datasets, outperforming SOTA methods on long videos and matching prior methods on short videos while using only one-eighth of the parameters.  [^1] The code is available at https://github.com/thearkaprava/MS-Temba
Figure 1: Temporal Action Detection using Mamba: On the left, a visual sample demonstrates that our model, MS-Temba, accurately localizes detected action labels in comparison to MS-TCT for temporal action detection in untrimmed videos. On the right, a plot of GPU memory usage versus video duration highlights the efficiency of MS-Temba in processing long untrimmed videos.
“Temba, his arms wide.”
Star Trek: The Next Generation episode,“Darmo”
@@ 1 Introduction
Learning representations in long, untrimmed videos is challenging due to complex, long-term dependencies that advanced neural networks struggle to optimize end-to-end. One such task, action detection, involves precisely localizing temporal boundaries of actions in untrimmed videos, with applications in areas like patient monitoring and video surveillance. Real-world deployment of temporal action detection models requires efficiency in capturing actions and their boundaries, yet real-world videos differ significantly from sports datasets [22, 2], where a few, non-overlapping actions with clear motion patterns are present. Real-world videos often contain densely distributed actions over time with subtle motion patterns, further complicated by large intra-class variance (e.g., the same action occurring with varying durations), the need for long-term relation modeling, and simultaneous actions. This paper focuses on developing an efficient model for action detection in hour-long videos, addressing these real-world challenges.
To learn temporal representations in untrimmed videos, early methods employed proposal-based approaches [43], which, while effective for videos with sparse action distributions, are resource-intensive due to the need to generate numerous proposals. Consequently, temporal convolutional methods [14, 31, 32] became popular for action detection, although their local receptive fields limit the capture of long-term dependencies. With the success of self-attention mechanisms, transformer-based methods [39, 8] were introduced to learn global representations across time frames. Despite their effectiveness, these transformer models impose high computational costs in terms of model parameters, memory usage, and throughput, making them challenging to deploy on edge devices for real-world applications.
In recent research, Mamba [17, 26, 50] is gaining attention for its efficiency in handling long sequences with linear complexity, making it well-suited for action detection in untrimmed videos. While Mamba has seen exploration in the vision domain, its application to action detection in untrimmed videos remains underexplored, motivating us to develop a Mamba-based architecture tailored for this purpose. To address this, we propose Multi-scale Temporal Mamba (MS-Temba), an architecture designed for efficient action detection (see Figure 1). MS-Temba incorporates two key components: 1) Temporal Mamba Blocks (Temba blocks) and 2) the Temporal Mamba Fuser (Temba fuser). The Temba Blocks are designed to learn temporal representations at multiple scales, achieved through two modules: the Temporal Local Module (TLM) and the Dilated Temporal State Space Module (DTS). TLM captures local temporal representations, while DTS learns long-range relationships using state space models enhanced with dilations. Although dilated convolutions are widely used, this work is the first to incorporate dilations into SSMs, allowing DTS to skip immediate neighbors and capture relationships across different temporal scales. This approach enhances Temba Blocks’ ability to model long-term temporal relations effectively. Additionally, the Temba Fuser aggregates scale-specific temporal information from each Temba Block, learning a multi-scale representation. This fusion, performed via Mamba, is computationally lighter than dense layers. We assess MS-Temba on three publicly available datasets with densely labeled actions, finding that it performs on-par with previous state-of-the-art method on short-duration videos and surpasses it on long-duration videos, using only one-eighth the parameters. Our study reveals that MS-Temba outperforms transformer-based models when both use a foundational visual backbone, highlighting Mamba’s potential not only for long sequences but also for feature inputs from foundation models [33].
To summarize the main contributions:
@@
    We introduce a novel Mamba-based architecture Multi-scale Temporal Mamba (MS-Temba), tailored for densely labeled Temporal Action Detection
@@
    MS-Temba explores different temporal scales through our novel components in Temba block, especially dilated Temporal SSMs, and aggregates them to learn multiscale representation through the Temporal Mamba Fuser
@@
    This carefully crafted design of MS-Temba  achieves performance on par with Transformer-based architectures on short videos and outperforms the previous state-of-the-art on long-duration videos with $88\%$ less parameters.
Figure 1 illustrates a visual sample of action detection using MS-Temba compared to previous state-of-the-art methods. MS-Temba efficiently processes hour-long videos with low GPU memory usage and fewer parameters, maintaining high accuracy. This motivates the suitability of the Mamba architecture for temporal action detection in untrimmed videos.
@@ 2 Related Work
Understanding actions in long temporal sequences has become a key task in computer vision [16, 27, 44, 11, 10, 48]. Early action detection approaches focused on videos with sparse action labels [22, 48, 2], primarily using proposal-based methods [43, 49] inspired by object detection to generate action proposals. However, temporal action detection with densely labeled action distribution offers greater challenges due to concurrent actions, where proposal-based methods struggle with the combinatorial explosion in proposal generation.
@@
This led to the adoption of temporal convolution-based models [31, 14] for processing long videos. Convolutional methods like PDAN [7] and TGM [32] use specialized kernels designed to model long-term dependencies and capture composite actions. However, the shared kernels in temporal convolutions are limited to neighboring frames, and even with dilations, struggle to capture global relationships between frames. This limitation is significant for densely labeled Temporal Action Detection, where long-term actions are common. In response, transformer-based architectures [41, 12, 40, 29, 13], which have proven effective in computer vision, have been adapted for action detection [39, 8, 34]. These architectures can model both short- and long-term dependencies in videos. For example, MLAD [39] employs a Transformer encoder to capture cross-class features and temporal relationships, while MS-TCT [8] uses a temporal hierarchy to model local relationships through temporal convolutions and global interactions through attention. These transformer based approaches enhances multi-scale representation learning, but the effectiveness comes at a high computational cost. The linear complexity of state space models (Mamba [17]) makes it an ideal test bed for processing long untrimmed videos.

### v2

@@ Abstract
Temporal Action Detection (TAD) in untrimmed videos requires models that can efficiently (1) process long-duration videos, (2) capture temporal variations within action classes, and (3) handle dense, overlapping actions, all while remaining suitable for resource-constrained edge deployment. While Transformer-based methods achieve high accuracy, their quadratic complexity hinders deployment in such scenarios. Given the recent popularity of linear complexity Mamba-based models, leveraging them for TAD is a natural choice. However, naïvely adapting Mamba from language or vision tasks fails to provide an optimal solution and does not address the challenges of long, untrimmed videos. Therefore, we propose Multi-Scale Temporal Mamba (MS-Temba), the first Mamba-based architecture specifically designed for densely labeled TAD tasks. MS-Temba features Temporal Mamba Blocks (Temba Blocks), consisting of Temporal Convolutional Module (TCM) and Dilated SSM (D-SSM). TCM captures short-term dependencies using dilated convolutions, while D-SSM introduces a novel dilated state-space mechanism to model long-range temporal relationships effectively at each temporal scale. These multi-scale representations are aggregated by Scale-Aware State Fuser, which learns a unified representation for detecting densely overlapping actions. Experiments show that MS-Temba achieves state-of-the-art performance on long-duration videos, remains competitive on shorter segments, and reduces model complexity by 88%. Its efficiency and effectiveness make MS-Temba well-suited for real-world edge deployment.  [^1] The code is available at https://github.com/thearkaprava/MS-Temba
Figure 1: Left: Temporal Action Detection poses unique challenges, including the need for computationally efficient models capable of processing long video sequences, the presence of both short and long actions with high intra-class variance, and the complexity of densely overlapping actions. Right: MS-Temba, our proposed model, achieves state-of-the-art performance while being $\mathbf{9\times}$ more parameter-efficient compared to transformer-based and convolution-based approaches. Furthermore, its lightweight architecture enables seamless deployment on edge devices, underscoring its efficiency and practicality for real-world applications.
@@ 1 Introduction
[…]
To learn temporal representations in untrimmed videos, early methods relied on proposal-based approaches [53], which, while effective for sparsely distributed actions, are computationally expensive due to dense proposal generation. To improve efficiency, Temporal Convolutional Networks (TCNs) [14, 39] have been explored, but their local receptive fields limit their ability to capture long-range dependencies. Recently, Transformer-based methods [49, 8] have achieved strong performance in TAD by leveraging global self-attention. However, their quadratic complexity leads to excessive memory consumption, making them impractical for long video sequences and edge deployment. Given the growing need for efficient long-sequence modeling, state-space models (SSMs) such as Mamba [18] have emerged, offering linear complexity while effectively capturing long-term dependencies. Although Mamba has demonstrated promising direction in vision tasks [18, 60, 27, 4], its application to TAD in untrimmed videos remains underexplored, motivating us to develop a Mamba-based architecture tailored for this purpose.
A naïve approach to adapting Mamba for TAD is to directly apply 1D Mamba to temporal segment representations, which are pooled frame-level features from a visual backbone. However, this approach sacrifices local information by compressing all context into a single state [18]. Additionally, Mamba sequentially processes tokens and relies on memorization, limiting its ability to model long-range temporal dependencies compared to Transformer counterparts. These limitations hinder its effectiveness in detecting actions in untrimmed videos. Therefore, we propose Multi-Scale Temporal Mamba (MS-Temba), a novel Mamba-based architecture designed for capturing temporal representations in videos. MS-Temba comprises two key components: (1) Temporal Mamba Blocks (Temba Blocks) and (2) Scale-Aware State Fuser (SSF), each tailored to address specific challenges in TAD. Temba Blocks capture both fine-grained and long-term dependencies by learning representations at multiple temporal scales. Each Temba Block consists of two modules: Temporal Convolutional Module (TCM) and Dilated SSM (D-SSM). TCM captures short-term dependencies using dilated convolutions, while D-SSM extends Mamba’s capability by incorporating dilations in the state-space update mechanism. The combination of TCM and D-SSM enables Mamba to learn discriminative temporal representations, with dilation parameters facilitating multi-scale temporal learning. To handle dense, overlapping actions spanning both short and long durations, effective aggregation of scale-aware temporal representations is essential. This is achieved through our proposed Scale-Aware State Fuser, which learns a unified representation across multiple temporal scales using SSMs.
Our experimental analysis shows that MS-Temba outperforms state-of-the-art performance on long-duration videos while maintaining comparable accuracy to Transformer-based models on short-duration videos, using only one-eighth of the parameters. Additionally, MS-Temba outperforms Transformer-based models when both utilize a foundational visual backbone like CLIP [40], demonstrating Mamba’s superior effectiveness in high-information-density settings.
While “local-global” representation learning and “multi-scale temporal learning” have been studied in prior work [14, 8, 35, 13, 28], MS-Temba is the first attempt to learn such representations leveraging SSMs for TAD. This paper presents a systematic approach to adapting the Mamba architecture for TAD, yielding key findings. Our carefully designed strategy enables MS-Temba to be trained directly on edge devices, providing advantages in privacy, personalization, efficiency, and real-time adaptability, making it well-suited for wearables and autonomous systems (see Figure 1). To summarize the main contributions:
@@
    We introduce Multi-Scale Temporal Mamba (MS-Temba), the first Mamba-based architecture tailored for densely labeled Temporal Action Detection.
@@
    We propose Temba Blocks, consisting of TCM and D-SSM. TCM captures short-term action dependencies using convolutions, while our novel D-SSM module introduces dilations in SSMs to model long-range temporal contexts. Additionally, we introduce Scale-Aware State Fuser, which efficiently aggregates multi-scale temporal representations, enabling dense action localization.
@@
    MS-Temba surpasses prior state-of-the-art methods on long-duration videos while maintaining competitive performance on shorter videos, with 88% fewer parameters.
@@ 2 Related Work
Figure 2: Overall Architecture of Multiscale Temporal Mamba (MS-Temba) for action detection. MS-Temba is composed of a frozen pretrained Visual Backbone, Temporal Mamba (Temba) Blocks for learning local-global representations at multiplle scale spaces. The Scale-Aware State Fuser employs an SSM for effectively fusing the multi-scale features which is projected by the Classification Head for dense action detection.
@@
Temporal Action Detection. Understanding actions in long temporal sequences has become a key task in computer vision [17, 30, 54, 11, 10, 58]. Early action detection approaches focused on videos with sparse action labels [23, 58, 2], primarily using proposal-based methods [53, 59] inspired by object detection to generate action proposals. However, temporal action detection with densely labeled action distribution offers greater challenges due to concurrent actions, where proposal-based methods struggle with the combinatorial explosion in proposal generation.
@@
This led to the adoption of temporal convolution-based models [38, 14, 31, 32] for processing long videos. Convolutional methods like PDAN [7] and TGM [39] use specialized kernels designed to model long-term dependencies and capture composite actions. However, the shared kernels in temporal convolutions are limited to neighboring frames, and even with dilations, struggle to capture global relationships between frames. This limitation is significant for densely labeled Temporal Action Detection, where long-term actions are common. In response, transformer-based architectures [51, 12, 50, 35, 13], which have proven effective in computer vision, have been adapted for action detection [49, 8, 43, 47, 16, 42]. These architectures can model both short- and long-term dependencies in videos. For example, MLAD [49] employs a Transformer encoder to capture cross-class features and temporal relationships, while MS-TCT [8] uses a temporal hierarchy to model local relationships through temporal convolutions and global interactions through attention. These transformer based approaches enhances multi-scale representation learning, but the effectiveness comes at a high computational cost. The linear complexity of state space models [18] makes it an ideal test bed for processing long untrimmed videos.